OpenStudy (anonymous):

Solve the system of linear equations using linear combination. 3a + 6b = 45 2a – 2b = –12 Which is the solution to the system?

OpenStudy (anonymous):

multiply the second equation all the way across by \(3\) then add the two equations in a column

OpenStudy (anonymous):

6a - 6b = -36

OpenStudy (anonymous):

a=15-2b

OpenStudy (anonymous):

a=b-6

OpenStudy (anonymous):

ok good you got \[3a + 6b = 45 \\2a – 2b = –12\] and turned it in to \[3a+6b=45\\ 6a-6b=-36\] now add them up

OpenStudy (anonymous):

what do you get when you add?

OpenStudy (anonymous):

9a = 9

OpenStudy (anonymous):

k good and that tells you \(a=1\)

OpenStudy (anonymous):

Its either two answers a=1, or its 9 or its -39